OpenStudy (anonymous):

Solve for x: 81^3-x=(1/3)^5x-1

OpenStudy (campbell_st):

use index laws Left hand side \[81^{3 - x} = (3^3)^{3 - x} = 3^{9 - 3x}\] Right hand side \[(1/3)^{5x -1 } = (3^{-1})^{5x - 1} = 3^{-5x + 1}\] now both sides are to the same base equate the powers 9 - 3x = -5x + 1 2x = -8 x = -4

OpenStudy (anonymous):

thanks!